OpenStudy (anonymous):

help me with trig please :( a boat sails 14 km west, then 12 km south. find its true bearing from its starting point.

OpenStudy (anonymous):

What is its starting point?

OpenStudy (anonymous):

that's all it says

OpenStudy (anonymous):

well I will make an example to hopefully explain to you how to solve it lets say the starting is (0,0)

OpenStudy (anonymous):

If it is 14 west and 12 south west being left south being down we will move 14 spaces left on the graph and then 12 spaces down and make a point there.

OpenStudy (anonymous):

I would double-check if I am correct with someone else as I am not 100% sure.

OpenStudy (phi):



OpenStudy (phi):

You should find the angle x. use trig: tan x = opposite/adjacent = 12/14 tan x = 6/7 \[ x = \tan^{-1}\left(\frac{6}{7}\right) \]